4 ELECTRIC FILED OF A DISK (20 POINTS) Two 10-cm diameter uniformly charged disks face each other, 20cm apart. The left disk is charged to 500 nC (1 nC 10-9 C) and the right disk is charged to +50nC. a) What is the electric field E (both magnitude and direction at the midpoint of the two disks? b) What electric force F (both magnitude and direction) on a 1.0 no charge placed at the midpoint?
Answer #1

a) from the given data
area of each disk, A = pi*d^2/4

= pi*0.1^2/4

= 7.85*10^-3 m^2

magnitude of surface charge density of each disk, sigma = Q/A

= 50*10^-9/(7.85*10^-3)

= 6.37*10^6 C/m^2

E = sigma/epsilon

= 6.37*10^-6/(8.854*10^-12)

= 7.19*10^5 N/c

direction : towards left disk

b) F = q*E

= 1*10^-9*7.19*10^5

= 7.19*10^-4 N

direction : towards right disk